OpenStudy (anonymous):

for this equation: ^3√24+^3√81-^3√54 is the answer ^3√51

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you can combine the third roots together and add and subtract the numbers under the roots

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

5(3^(1/3))-3(2^(1/3))

OpenStudy (anonymous):

\[5\sqrt[3]{3}-3\sqrt[3]{2}\]

OpenStudy (anonymous):

it gives me both of does answers

OpenStudy (anonymous):

pick the second one, it is not \[\sqrt[3]{51}\]

OpenStudy (anonymous):

ok